OpenStudy (anonymous):

quick copy recently charged 49 cents per page for color copies and 7 cent per page for black and white copies. if shirly's bill for 90 copies was 11.34 how many copies of eahc type were made.

OpenStudy (anonymous):

c is color copies b is black and white copies 49c + 7b = 1134 c + b = 90 ---> c = 90 - b Now substitute 90 - b in for c in the first equation... 49c + 7b = 1134 49(90 - b) + 7b = 1134 4410 - 49b + 7b = 1134 -49b + 7b = 1134 - 4410 - 42b = - 3276 b = 3276/42 b = 78 Now substitute 78 in for b in the second equation... c = 90 - b c = 90 - 78 c = 12 check...remember c = 12 and b = 78 .49c + .07b = 11.34 .49(12) + .07(78) = 11.34 5.88 + 5.46 = 11.34 11.34 = 11.34 (correct) ANSWER : Shirly bought 12 color copies at .49 cents each, totaling $ 5.88 and 78 black and white copies at .07 cents each, totaling $ 5.46. Total is 90 copies for $ 11.34.